OPTICAL ZEEMAN SPECTROSCOPY OF IRON MONOHYDRIDE, FeH

JINHAI CHEN, TIMOTHY C. STEIMLE^{a}, Department of Chemistry and Biochemistry, Arizona State University, Temple, AZ 85287; \underline{ZHONG WANG}, TREVOR J. SEARS^{b}, Department of Chemistry, Brookhaven National Laboratory, Upton, NY 11973-5000.

The polarization signals induced by the Zeeman effect in the X^{4}Δ ← F^{4}Δ band of FeH have been proposed for use in the characterization of stellar magnetic fields.^{c} While the magnetic tuning of the low-rotational levels of the X^{4}Δ_{7/2} state has been measured by pure-rotational laser magnet resonance spectroscopy,^{d} the strength of the Zeeman effect in the upper state is unknown. Here we report on the magnetic tuning of levels in (v=1)F^{4}Δ_{7/2} state derived from the analysis of the optical Zeeman spectra of the X^{4}Δ_{7/2} ← F^{4}Δ_{7/2} (1,0) band system. The Zeeman broadening and splitting of the Q(3.5)(ν = 11356.72 cm^{−1}) transition were successfully modeled using a traditional effective Hamiltonian. The determined magnetic g_{L}-factor is 1.079(8) when the g_{S}-factor is constrained to 2.002. Plausible explanations for deviation of g_{L} from the expected value of unity will be given. The determined parameters are used to predict the magnetic tuning of the X^{4}Δ_{7/2} ← F^{4}Δ_{7/2} (1,0) features for fields comparable to those found in sunspots.

aThis research was supported by the National Science Foundation, Experimental Physical Chemistry Division.

bWork at Brookhaven National Laboratory was carried out under Contract No. DE-AC02-98CH10886 with the U.S. Department of Energy and supported by its Division of Chemical Sciences, Office of Basic Energy Sciences.

cS.V.Berdyugina, and S.K.Solanki, Astron. Astrophys. 385, 201(2002)

dJ.M.Brown, H.Korsgen, S.P.Beaton, and K.M.Evenson (Submitted to J.Chem.Phys.)